# Patterns of Non-Stationary Solutions to Symmetric Systems of Second-Order Differential Equations

### v1

@@ Appendix A The $G$-Equivariant Leray-Schauder Degree
to specify the integer coefficient standing next to the generator element $(H)\in\Phi_{0}(G)$. An Axiomatic Construction of the Leray-Schauder $G$-Equivariant Degree. Let $\mathscr{E}$ be any isometric Banach $G$-representation. A map $f:\mathscr{E}\rightarrow\mathscr{E}$ is said to be a completely continuous $G$-equivariant field if it can be expressed in the form
@@
Computation of Brouwer equivariant degree. Let’s denote by $\{\mathcal{U}_{i}\}_{i\in\mathbb{N}}$ the set of all irreducible $G$-representations and define the $i$-th basic degree as follows
@@
and any bounded $G$-equivariant linear isomorphism $T:\mathscr{E}\to\mathscr{E}$, the Multiplicativity and Homotopy properties of the $G$-equivariant Brouwer degree, together with Schur’s Lemma implies

### v2

@@ Appendix A The $G$-Equivariant Leray-Schauder Degree
to specify the integer coefficient standing next to the generator element $(H)\in\Phi_{0}(G)$. An Axiomatic Construction of the $G$-equivariant Leray-Schauder-Equivariant degree. Let $\mathscr{E}$ be any isometric Banach $G$-representation. A map $f:\mathscr{E}\rightarrow\mathscr{E}$ is said to be a completely continuous $G$-equivariant field if it can be expressed in the form
@@
Computation of the $G$-equivariant Leray-Schauder degree. Let’s denote by $\{\mathcal{U}_{i}\}_{i\in\mathbb{N}}$ the set of all irreducible $G$-representations and define the $i$-th basic degree as follows
@@
and any bounded $G$-equivariant linear isomorphism $T:\mathscr{E}\to\mathscr{E}$, the Multiplicativity and Homotopy properties of the $G$-equivariant Leray-Schauder degree, together with Schur’s Lemma implies